OpenStudy (anonymous):

Still a little confused on what equation to use and how to figure this out! Any one please help me out? Find the slope of a line that passes through the points (2, –3) and (–1, 4).

OpenStudy (nincompoop):

first solve for the slope then use the y-intercept form to get the equation

OpenStudy (nincompoop):

do you know how to solve for the slope @SleepEatRugby

OpenStudy (the_fizicx99):

M = y2 - y1/X2 - x1

OpenStudy (sweetburger):

use point slope

OpenStudy (nincompoop):

@sweetburger @tHe_FiZiCx99 I got this... you two can calm down

OpenStudy (sweetburger):

point slope form i tell you

OpenStudy (the_fizicx99):

@nincompoop oki :(

OpenStudy (nincompoop):

@SleepEatRugby you need to respond

OpenStudy (anonymous):

I keep getting a negative number as the denominator for the slope would this be wrong?

OpenStudy (nincompoop):

show me how you are solving for the slope

OpenStudy (anonymous):

(4 - -3)/2 + (-1 - 2)/2

OpenStudy (nincompoop):

\[slope=m=\frac{ rise }{ run }=\frac{ \Delta y }{ \Delta x }=\frac{ y_2-y_1 }{ x_2-x_1 }=\frac{ 4-(-3) }{ -1-2 }\]

OpenStudy (nincompoop):

where are you getting the denominator 2? I think your mixing up mid-point and slope formulas

OpenStudy (nincompoop):

then whatever your m (slope) is plug it into the point-slope form \[y-y_1=m(x-x_1)\] use either (2, -3) or (-1, -4) as your x1 and y1

OpenStudy (nincompoop):

are you still there?

OpenStudy (the_fizicx99):

Need help?